OpenStudy (anonymous):

Find the equation of the plane through (1,1,1) and containing the line which is the intersection of the planes x-y=2 and y-z=1

OpenStudy (anonymous):

This is what i have tried doing:I got the vectors perpendicular to the gives planes, namely: v=i-j and u=j-k, then i took the cross product. This resulant vector is parrallel to the line of intersection. Now i believe i am in a position to write the equation of the plane since i have a vector perpinducalr to the plane and a point

OpenStudy (anonymous):

Equaition of plane through intersection of x-y-2=0;y-z-1=0 is (x-y-2)+a(y-z-1)=0 Put in this (1,1,1) and find a

OpenStudy (anonymous):

what would be the "a"

OpenStudy (anonymous):

a=-2

OpenStudy (anonymous):

okay, then what would be the final eequation of the plane

OpenStudy (anonymous):

i dont think that is correct

OpenStudy (anonymous):

NO it is correct

OpenStudy (anonymous):

okay then what is the final equation of the plane

OpenStudy (anonymous):

x-3y+2z=0